# Robust Continuous-Time Beam Tracking with Liquid Neural Network

### v1

@@ I Introduction
Millimeter-wave (mmWave) communication is a promising enabling technology for improving the data rate of wireless communications due to its abundance of spectrum resources at high frequencies [1, 2]. To fully exploit this potential, massive multiple-input multiple-output (MIMO) configurations have been deployed in both base stations (BSs) and user equipment (UE) to increase system throughput. However, the high-dimensional signal processing associated with massive MIMO imposes a significant computational overhead [3]. Additionally, the presence of background noise in urban environments with complex electromagnetic landscapes poses a substantial challenge in determining the optimal beam directions.
@@
Machine learning (ML) is adept at leveraging historical data to address new problems with similar data structures. Previous researches have demonstrated that ML-based techniques can effectively perform beamforming with mmWave [5, 6, 7, 8, 9, 10, 11]. Specifically, long short-term memory (LSTM) neural networks have been employed to predict beams based on past pilot signals [8, 9]. Through nonlinear fitting to model the features of UE movements, the variations in AoA and AoD can be more precisely captured, resulting in improved performance compared to conventional methods. However, due to the inherent design of LSTM for discrete-time processing, it falls short in real-time beam adjustments [10]. To overcome this limitation, [11] integrated the LSTM structure with ordinary differential equation (ODE) to predict the optimal beam at arbitrary instants. By incorporating ODEs to model the derivative of beam variations, this approach achieved higher beamforming gains in continuous-time beam tracking tasks.
Nonetheless, the above approaches overlooked the deployment of beam tracking in urban areas characterized by intricate electromagnetic landscapes, where users are frequently moving amidst substantial interference noise. Traditional deep learning methods exhibited limited robustness in such noisy settings, leading to a significant degradation in performance under heavy interference. To tackle the challenges posed by dynamic scenarios with noise, a specialized type of recurrent neural network called liquid neural network (LNN) has been introduced [12]. Inspired by the nervous system of Caenorhabditis elegans, LNN emulates the time-varying behavior of synapses in the brain, which enables the network to excel at continuous-time data processing. It demonstrated exceptional adaptability in various applications such as autopilot systems [13] and flight navigation [14]. In this paper, we propose a robust continuous-time beam tracking scheme employing LNN. Specifically, the received pilot signal vectors undergo feature extraction, then the extracted features are processed by the LNN unit. Finally, the output layer produces a probability vector that determines the selection of the beam from the codebook. Leveraging the modeling of LNN, our proposed method can predict the sub-optimal beam at arbitrary instants. Furthermore, the flexibility of the nonlinear response in the designed model enables our scheme to adaptively process rapidly changing noisy channels. Simulation results confirm the effectiveness of our proposed method in achieving up to 46.9% higher normalized spectral efficiency than the state-of-the-art baselines with the UE moving at 5 m/s.

### v2

@@ I Introduction
Millimeter-wave (mmWave) communication is a promising enabling technology for improving the data rate of wireless communications due to its abundance of spectrum resources at high frequencies [1, 2, 3]. To fully exploit this potential, massive multiple-input multiple-output (MIMO) configurations have been deployed in both base stations (BSs) and user equipment (UE) to increase system throughput [4]. However, the high-dimensional signal processing associated with massive MIMO imposes a significant computational overhead [5]. Additionally, the presence of background noise in urban environments with complex electromagnetic landscapes poses a substantial challenge in determining the optimal beam directions [6].
@@
Machine learning (ML) is adept at leveraging historical data to address new problems with similar data structures [8, 9]. Previous researches have demonstrated that ML-based techniques can effectively perform beamforming with mmWave [10, 11, 12, 13]. Specifically, long short-term memory (LSTM) neural networks have been employed to predict beams based on past pilot signals [14, 15]. Through nonlinear fitting to model the features of UE movements, the variations in AoA and AoD can be more precisely captured, resulting in improved performance compared to conventional methods. However, due to the inherent design of LSTM for discrete-time processing, it falls short in real-time beam adjustments [16]. To overcome this limitation, [17] integrated the LSTM structure with ordinary differential equation (ODE) to predict the optimal beam at arbitrary instants. By incorporating ODEs to model the derivative of beam variations, this approach achieved higher beamforming gains in continuous-time beam tracking tasks.
Nonetheless, the above approaches overlooked the deployment of beam tracking in urban areas characterized by intricate electromagnetic landscapes, where users are frequently moving amidst substantial interference noise. Traditional deep learning methods exhibited limited robustness in such noisy settings, leading to a significant degradation in performance under heavy interference [18]. To tackle the challenges posed by dynamic scenarios with noise, a specialized type of recurrent neural network called liquid neural network (LNN) has been introduced [19]. Inspired by the nervous system of Caenorhabditis elegans, LNN emulates the time-varying behavior of synapses in the brain, which enables the network to excel at continuous-time data processing. It demonstrated exceptional adaptability in various applications such as autopilot systems [20] and flight navigation [21]. In this paper, we propose a robust continuous-time beam tracking scheme employing LNN. Specifically, the received pilot signal vectors undergo feature extraction, then the extracted features are processed by the LNN unit. Finally, the output layer produces a probability vector that determines the selection of the beam from the codebook. Leveraging the modeling of LNN, our proposed method can predict the sub-optimal beam at arbitrary instants. Furthermore, the flexibility of the nonlinear response in the designed model enables our scheme to adaptively process rapidly changing noisy channels. Simulation results confirm the effectiveness of our proposed method in achieving up to 46.9% higher normalized spectral efficiency than the state-of-the-art baselines with the UE moving at 5 m/s.
The rest of this paper is organized as follows: Section II describes the system model and problem formulation. Section III provides a comprehensive analysis of the LNN-based beam tracking method, including the architecture of LNN and the flowchart of the proposed deep learning model. Simulation results are presented in Section IV to verify the performance of the proposed algorithm. Finally, the conclusions are outlined in Section V.
@@ V Conclusion
Acknowledgement
The work was supported by the China National Key R&D Program under Grant 2021YFA1000500 and 2023YFB2904804, National Natural Science Foundation of China under Grant 62331023, 62101492, 62394292 and U20A20158, Zhejiang Provincial Natural Science Foundation of China under Grant LR22F010002, Zhejiang Provincial Science and Technology Plan Project under Grant 2024C01033, and Zhejiang University Global Partnership Fund, MOE Tier 2 (Award number MOE-T2EP50220-0019) and A*STAR (Agency for Science, Technology and Research) Singapore, under Grant No. M22L1b0110.